The annual interest on an $18,000 investment exceeds the interest earned on an
$8000 investment by $290.
The $18,000 is invested at a 0.5% higher rate of interest than the $8000.
What is the interest rate of each investment?

Answer by stanbon(75887)

The annual interest on an $18,000 investment exceeds the interest earned on an
$8000 investment by $290.
----
18000(x/100) = 8000(y/100) + 290
------------------------------------
The $18,000 is invested at a 0.5% higher rate of interest than the $8000.
What is the interest rate of each investment?
x = y + 0.5
----------------
Substitute for "x" and solve for "y":
180(y+0.5) = 80y + 290
-----
100y + 90 = 290
-----
100y = 200
y = 2% (interest on the $18000)
x = y+0.5 = 2.5% (interest on the $8000
--------
Cheers,
Stan H.
--------------